OpenStudy (anonymous):

Evaluate the following expressions. The answer must be given as a fraction, NO DECIMALS. If tan=−3/4 and cosX>0, then sinx= cosx= cotx= secx= cscx= x is acting as theta.

OpenStudy (anonymous):

@RadEn help?

OpenStudy (raden):

i sure x in th4 quadrant

OpenStudy (anonymous):

idk... i can do it with out restrictions... but with i get totally thrown off...

OpenStudy (raden):



OpenStudy (anonymous):

5

OpenStudy (raden):

smart :)

OpenStudy (anonymous):

thanks i try ^_^

OpenStudy (anonymous):

so sin should be -3/5... but that is less than one...

OpenStudy (raden):

yup, that's correct dont worry , intervals for sin -1<= sin <= 1

OpenStudy (anonymous):

awww snap (: thanks a bundle ! so, cos= 4/5 sec= -5/3 csc= 5/4 cot=-4/3

OpenStudy (raden):

sec just follows with cos, if cos positive then sec positive also

OpenStudy (anonymous):

but not in that quadrant though?

OpenStudy (raden):